OpenStudy (anonymous):

At the same moment from the top of a building 3.0 × 10 2 m tall, one rock is dropped and one is thrown downward with an initial velocity of 23 m/s. Both of them experience negligible air resistance. How much EARLIER does the thrown rock strike the ground?

OpenStudy (anonymous):

multiply then divide?

OpenStudy (anonymous):

multiply what and divide what

OpenStudy (anonymous):

use the speed formula.

OpenStudy (anonymous):

3.0 × 10 2

OpenStudy (anonymous):

V = d/t i suppose you are required to find time. rearrange equation to find time.

OpenStudy (anonymous):

how i did 3.0*10^2/23 but the answer is not correct

OpenStudy (anonymous):

where u go

OpenStudy (igreen):

https://answers.yahoo.com/question/index?qid=20130609031821AAsLSHq You need to use the equation: s = ut + (!/2)at^2 you can figure out the time that the dropped ball to hit the ground quite easily because u = 0 m/s, therefore s = (1/2)at^2 t^2 =2s/a t = sqrt(2s/a) assuming the acceleration from gravity is 9.81 m/s^2 t = sqrt(2*300m/9.81 m/s*2) t = 7.8 s You can use the quadratic equation to find t for the thrown rock and then subtract them. Or you can calculate the final velocity of the thrown rock with v^2 = u^2 + 2as and then use v = u + at to find the time for the thrown rock. v^2 = (10 m/a)^2 + 2 * 9.81 m/s^2 * 300 m v = 76.7 m/s v = u + at t = (v - u)/s t = (76.7 m/s - 10 m/s)/9.81 m/s^2 t = 6.8 sec So the throw rock hit the ground one second earlier than the dropped rock.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @gamal992

OpenStudy (alekos):

isn't u = 23m/s?

OpenStudy (anonymous):

i know throw rock hit the ground earlier but there are a choices -1.4 s -1.8 s -2 s